OpenStudy (anonymous):

Sinx-2sin^2 x=0 Step by step?

OpenStudy (blurbendy):

you want to solve for x?

OpenStudy (mrhoola):

yes that is what she needs

OpenStudy (anonymous):

Yes please!

OpenStudy (blurbendy):

ok

OpenStudy (blurbendy):

first factor our sinx and the constants -(sin(x)(2sin(x) -1) = 0 Multiply both sides by a constant to simplify sin(x)(2sin(x) -1) = 0 solve each term in the product separately sin(x) = 0 2sin(x) - 1 = 0 Can you proceed from here?

OpenStudy (anonymous):

Yes thank you so much!